# Causally constrained reduced-order neural models of complex turbulent dynamical systems

### v1

@@ II Framework: from linear response to causal constraints
The proposed constraints leverage information about zeros of the ensemble-averaged Jacobian to constrain the state-dependent Jacobian. While this choice is practical and efficient, it can be limiting in idealized systems, for example because of exact symmetries. In such cases, nonzero local couplings may cancel in ensemble averages, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$ even if $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$, as occurs, for example, in the Lorenz ’63 model [58]. In such idealized settings, applying the proposed approach may require a symmetry-reduction step, as advocated in [22] and demonstrated in [77] for the Lorenz ’63 system. These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations. In this regime, dominant interactions among coarse-grained variables are often well approximated by a state-independent linear operator, with nonlinearities providing higher-order corrections [88, 87, 89, 90, 111, 41, 64, 70]. There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics. Further details are provided in Section II of the SM.
@@ II Possible limitations of the causal constraints and practical considerations
Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [29]. Specifically, the response operator at the shortest time-scale can be written in terms of the ensemble averaged Jacobian as $\mathbf{R}(dt)=\langle\mathbf{J}(\mathbf{x})\rangle dt+\mathbf{I}$, where $\mathbf{J}(\mathbf{x})$ denotes the Jacobian, $\mathbf{I}$ is the identity matrix, and $dt\rightarrow 0$. Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss. In other words, the central assumption of this work is that knowledge of the zeros of the ensemble averaged Jacobian already provides an informative constraint for guiding the architecture of general neural emulators. A complication arises in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [58, 59]. In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise. In such cases, a vanishing short-time response does not necessarily imply the absence of an instantaneous coupling. One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches as advocated in [22]. For example, in the Lorenz ’63 system [58], this corresponds to considering the so-called “proto-Lorenz” system [77], in which the symmetries of the original system are reduced. More importantly, however, we argue that this limitation does not constitute a practical obstacle for reduced-order modeling of real-world turbulent climate systems, which are the primary focus of this work. Such systems typically lack exact symmetries and admit no unique coarse-grained representation. The effective drift term $\mathbf{f}$ is validated on its ability to reproduce stationary and statistics given a specific choice of coarse-grained representation. In this context, we find it useful to provide additional qualitative justification for why focusing on zeros of the ensemble-averaged Jacobian is not a fundamental limitation. First, we discuss a qualitative mathematical argument in the context of linear inverse models of climate. Second, we emphasize a more general point regarding the non-uniqueness of effective dynamics in data-driven modeling of coarse-grained systems.
  • In the context of reduced-order modeling of large-scale climate variability, typically constructed from anomalies, there is extensive evidence that linear inverse models (LIMs), i.e. $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W}$, as originally proposed by Penland [88], provide a useful effective approximation of the dynamics and serve as powerful forecasting tools; see [88, 87, 89, 90, 111, 41] among many others. A natural generalization incorporates quadratic nonlinearities,
    $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x}))dt+\mathbf{\Sigma}d\mathbf{W}$
    as discussed, for example, in [64]. In what follows, we focus on nonlinear extensions of the LIM within this quadratic framework. In this case, the $k$-th component of the drift can be written as
    $f_{k}(\mathbf{x})=\sum_{m}L_{k,m}x_{m}+\sum_{m,n}B_{k,m,n}x_{m}x_{n}~.$
    and the corresponding Jacobian entry takes the form
    $J_{k,j}(\mathbf{x})=\partial f_{k}(\mathbf{x})/\partial x_{j}=L_{k,j}+\sum_{m}B_{k,j,m}x_{m}+\sum_{n}B_{k,n,j}x_{n}~.$
    Taking the ensemble average yields
    $\langle J_{k,j}(\mathbf{x})\rangle=\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=L_{k,j}+\sum_{m}B_{k,j,m}\langle x_{m}\rangle+\sum_{n}B_{k,n,j}\langle x_{n}\rangle~.$
    In the common modeling regime where climate variability is represented in terms of anomalies, one typically has $\langle x_{n}\rangle=0$ for all components, leading to
    $\langle J_{k,j}(\mathbf{x})\rangle=L_{k,j}~.$
    In this modeling setting, the ensemble-averaged Jacobian therefore coincides with the linear operator $\mathbf{L}$. The empirical success of linear inverse models for coarse-grained climate variability suggests that, at this level of coarse-grained description (typically around 20 principal components of a given climate field, see e.g. [87]), the linear operator $\mathbf{L}$ approximates the dominant direct couplings, with nonlinear terms acting as corrections. This linear operator is obviously state independent. As a result, in this example, any vanishing entry $\langle J_{k,j}\rangle=L_{k,j}=0$ reflects the absence of a linear coupling, rather than a cancellation induced by ensemble averaging, therefore providing intuition for why vanishing entries of the ensemble-averaged Jacobian are structurally meaningful constraints.
@@
  • More broadly, this work highlights a principle relevant to data-driven modeling of realistic multiscale systems. As discussed in [29], the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations. In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse graining. From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
@@
Finally, causally constrained emulators can be readily compared with an unconstrained counterpart by evaluating their representation of the linear response operator against the FDT-based estimate in Eq. (1) of the main text. Such comparisons should be performed at short time lags in order to avoid spurious responses that may arise at longer lags when using the FDT [28]. The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model. In systems with exact symmetries, as discussed above, any biases introduced by the constraints will already be apparent in the representation of stationary variability.

### v2

@@ II Framework: from linear response to causal constraints
The proposed constraints leverage information about zeros of the ensemble-averaged Jacobian to constrain the state-dependent Jacobian. While this choice is practical and efficient, it can be limiting in idealized systems, for example because of exact symmetries. In such cases, nonzero local couplings may cancel in ensemble averages, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$ even if $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$, as occurs, for example, in the Lorenz ’63 model [57]. In such idealized settings, applying the proposed approach may require a symmetry-reduction step, as advocated in [22] and demonstrated in [76] for the Lorenz ’63 system. These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations. There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics. Further details are provided in Section II of the SM.
@@ II Possible limitations of the causal constraints and practical considerations
Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [29]. Specifically, the response operator at the shortest time-scale can be written in terms of the ensemble averaged Jacobian as $\mathbf{R}(dt)=\langle\mathbf{J}(\mathbf{x})\rangle dt+\mathbf{I}$, where $\mathbf{J}(\mathbf{x})$ denotes the Jacobian, $\mathbf{I}$ is the identity matrix, and $dt\rightarrow 0$. Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss.
A complication can in principle arise in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [57, 58]. In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise. One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches advocated in [22]. For example, Ref. [76] proposed a nonlinear coordinate transformation to quotient the $Z_{2}$ symmetry of the Lorenz ’63 system [57], leading to the so-called “proto-Lorenz” system. In the case of the Lorenz ’96 system, defined as $\dot{u_{j}}=(u_{j+1}-u_{j-2})u_{j-1}-u_{j}+F,~j=0,...,J$ with periodic boundary conditions, the model is translationally invariant, with each $u_{j}$ having the same ensemble mean $\langle u_{j}\rangle=\bar{u}$. In this latter case, introducing a site-dependent forcing $F_{j}$ is sufficient to guarantee the validity of our framework. For example, one could define $F_{j}=F+\eta_{j}$, with $\eta_{j}\sim\mathcal{N}(0,\sigma^{2})$, mimicking the effect of topography, and therefore leading to a more realistic forcing term.
@@
Crucially, we argue that the limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems, with no symmetries. As discussed in [29], in the case of climate modeling, the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations. In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse-graining. From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
@@
Finally, causally constrained emulators can be readily compared with an unconstrained counterpart by evaluating their representation of the linear response operator against the FDT-based estimate in Eq. (1) of the main text. Such comparisons should be performed at short time lags to avoid spurious responses that may arise at longer lags when using the FDT [28]. The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model. In systems with exact symmetries, as discussed above, any biases introduced by the constraints will be immediately apparent in the representation of stationary variability.